1. Break-Even Sales and Sales Mix for a Service Company Zero Turbulence Airline provides air transportation...

1.

Break-Even Sales and Sales Mix for a Service Company

Zero Turbulence Airline provides air transportation services between Los Angeles, California; and Kona, Hawaii. A single Los Angeles to Kona round-trip flight has the following operating statistics:

Fuel $8,883
Flight crew salaries 6,804
Airplane depreciation 3,213
Variable cost per passenger—business class 50
Variable cost per passenger—economy class 40
Round-trip ticket price—business class 520
Round-trip ticket price—economy class 260

It is assumed that the fuel, crew salaries, and airplane depreciation are fixed, regardless of the number of seats sold for the round-trip flight. If required round the answers to nearest whole number.

a. Compute the break-even number of seats sold on a single round-trip flight for the overall product, E. Assume that the overall product is 20% business class and 80% economy class seats.

Total number of seats at break-even seats

b. How many business class and economy class seats would be sold at the break-even point?

Business class seats at break-even seats
Economy class seats at break-even seats

2.

Margin of Safety

a. If Canace Company, with a break-even point at $409,600 of sales, has actual sales of $640,000, what is the margin of safety expressed (1) in dollars and (2) as a percentage of sales? Round the percentage to the nearest whole number.

1. $

2.   %

b. If the margin of safety for Canace Company was 25%, fixed costs were $1,308,750, and variable costs were 75% of sales, what was the amount of actual sales (dollars)?
(Hint: Determine the break-even in sales dollars first.)
$
Answer #1
1
a. Total number of seats at break-even=Fixed costs/Weighted average contribution margin per passenger
Fixed costs:
$
Fuel $8,883
Flight crew salaries 6,804
Airplane depreciation 3,213
Total $18,900
Weighted average contribution margin per passenger:
Class Round-trip ticket price Variable cost per passenger Contribution margin per passenger Weight (Sales mix) Weighted average contribution margin per passenger
a b c=a-b d e=c*d
Business class 520 50 470 20% 94
Economy class 260 40 220 80% 176
Total 270
Total number of seats at break-even=18900/270=70 seats
b. Business class seats at break-even=Total number of seats at break-even*Sales mix %=70*20%=14 seats
Economy class seats at break-even=Total number of seats at break-even*Sales mix %=70*80%=56 seats
2
a. Margin of safety expressed in dollars=Actual sales-Break even point sales=640000-409600=$ 230400
Margin of safety expressed as a % of sales=Margin of safety expressed in dollars/Actual sales=230400/640000=0.36=36%
b. Beak-even in sales dollars=Fixed costs/Contribution margin %
Contribution margin %=100%-Variable cost %=100-75=25%
Beak-even in sales dollars=1308750/25%=$ 5235000
Margin of safety expressed as a % of sales=Margin of safety expressed in dollars/Actual sales
Margin of safety expressed in dollars=Actual sales-Break even point sales
Margin of safety expressed as a % of sales=(Actual sales-Break even point sales)/Actual sales
25%=(Actual sales-5235000)/Actual sales
25%*Actual sales=Actual sales-5235000
0.25Actual sales=Actual sales-5235000
Actual sales-0.25Actual sales=5235000
0.75Actual sales=5235000
Actual sales=5235000/0.75=$ 6980000